# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ I Introduction
The Tor network is a decentralized, volunteer-operated system designed to provide website browsing anonymity. Tor applies sophisticated onion routing through a series of relay nodes and encryption mechanisms to anonymize traffic. Its architecture is built to protect users from surveillance, traffic analysis, and censorship by hiding their identity and online activity. Thus, Tor makes adversaries difficult to trace the origin or destination of the data [1, 2, 3, 4].
However, Tor’s network traffic can still be susceptible to traffic analysis techniques, including website fingerprinting (WF) attacks [5, 6, 7]. By exploiting patterns in encrypted traffic, these attacks aim to expose the websites a user is visiting. Distinctive patterns in packet size [8, 9, 10], timing [11, 12, 13], and flow direction [12, 14] can still leak identifying information about the websites being accessed. In response, a variety of defense mechanisms [15, 16, 17] have been developed to obscure or randomize traffic patterns, thereby reducing the risk of identification. These strategies include traffic obfuscation [18, 19], adversarial perturbation [20, 21], and website morphing [22, 23, 24]. While effective in mitigating fingerprinting attempts, such defenses must carefully balance privacy protection with performance to ensure they do not degrade the user experience.
@@ I-A Motivation
Despite the growing significance of WF attack and defense approaches, a systematic and comprehensive survey of the field remains lacking in academia. Existing surveys primarily focus on WF attacks and defenses using machine learning (ML) models [25, 26, 27]. However, these works provide limited insights into the design of models in critical aspects such as concept drift, single-tab versus multi-tab browsing scenarios, and dynamic adversarial capabilities. Moreover, while previous surveys [27, 28] introduce several defense mechanisms, their taxonomies are neither systematic nor sufficiently granular. For instance, Chao et al. [28] categorize WF attacks and defenses within anonymity networks based on different protocol layers of the onion route. However, a gap remains in understanding the comparative effectiveness of various techniques under diverse threat models and defense scenarios. Existing work fails to adequately categorize defenses according to design principles (e.g., obfuscation, padding, traffic morphing) or threat model alignment, limiting their utility for comparative evaluation. Furthermore, datasets in shaping WF research are not well justified. These surveys fail to categorize defenses based on factors such as design principle (e.g., obfuscation, padding, traffic morphing) or threat model alignment adequately. For example, the existing survey [27] discusses dataset features and assumptions, but does not examine the availability, diversity, and limitations of existing datasets in depth. Therefore, a systematic review of WF attacks and defenses in Tor is imperative. Unlike existing surveys, our work provides an in-depth review of available datasets and examines their features to guide dataset selection for future research. Building on this foundation, we present a comprehensive review of WF attack approaches and categorize defense strategies according to their underlying mechanisms, highlighting their respective strengths and weaknesses against state-of-the-art (SOTA) attack models.
@@ I-B1 Datasets
Labeled datasets mapping traffic traces to websites are required in the SOTA WF attacks [29, 30, 31, 32, 33], which usually employ supervised learning techniques. The effectiveness of existing ML models largely depends on the quality, diversity, and scale of the datasets used. Therefore, the development of realistic, diverse, and up-to-date datasets is essential for advancing research in this field and ensuring the reliability of experimental results [34].
There are two types of datasets: closed-world and open-world. A closed-world dataset assumes that the user visits only a predefined set of websites (e.g., the top 100 sites from Alexa), and both the attacker and defender know this set in advance [35]. The closed-world dataset is better suited for studying the upper bound of attack performance in a controlled setting, as it facilitates comparisons between different WF models under ideal conditions. However, closed-world evaluation overlooks the fact that users can visit websites outside the training set of an ML model. In contrast, an open-world dataset contains a smaller set of monitored websites (the attacker is interested in) and a much larger set of unmonitored websites representing background traffic. In this setting, the attacker must distinguish between monitored and unmonitored sites. Open-world datasets more accurately reflect real-world usage, allowing researchers to evaluate the effectiveness of attacks and defenses under more realistic assumptions [9, 36].
@@ I-B2 WF attacks
The WF attacks section introduces the existing website fingerprinting attacks by providing a comprehensive overview of the various techniques that have been developed over time. This includes a detailed discussion of the underlying models used, ranging from traditional ML algorithms such as k-Nearest Neighbors (k-NN) [9, 33, 35] and Support Vector Machines (SVM) [35, 37, 4] to more advanced DL approaches such as Convolutional Neural Networks (CNN) [38, 39, 40, 13, 41, 42] and Transformer-based architectures [43, 30, 44]. The section also covers the process of feature selection, highlighting which traffic characteristics, such as packet sizes, directions, timing, and burst patterns, are most informative for classification. Additionally, the performance of different attacks is compared using standard evaluation metrics, such as accuracy, precision, and recall, across both closed-world and open-world scenarios. This comparison can help illustrate the evolution of WF attacks and provide insight into which methods are most effective under various conditions.
@@ I-B3 WF defenses
The WF defenses section introduces the existing techniques designed to counter website fingerprinting attacks, with a particular focus on a range of key defense mechanisms. Among these, adaptive padding is a dynamic approach that injects dummy traffic based on live traffic patterns to conceal burst and packet size features, striking a balance between privacy and bandwidth overhead. Traffic morphing attempts to transform the traffic of one website to resemble that of another, thereby confusing fingerprinting classifiers by altering key statistical features [9, 45, 23, 24]. Adversarial perturbation techniques borrow principles from adversarial ML to proactively disrupt attack models [46, 47, 48], where minor, intentional modifications to network traffic are used to mislead ML models. Each method is discussed in terms of its effectiveness, overhead, implementation complexity, and suitability for real-world Tor systems. This in-depth review will help readers understand the current state of WF defenses and the challenges that remain in developing robust, low-cost, and scalable solutions.
@@ II Related Works
Basyoni et al. [49] 2020 Traffic Analysis Attacks on Tor.
Karunanayake et al. [50] 2021 De-Anonymisation Attacks on Tor.
Aminuddin et al. [27] 2023 Analysis on Tor-based WF techniques and assumptions.
Liu et al. [25] 2023 Deep learning on Tor-based WF attacks and defenses.
Chao et al. [28] 2024
@@
Basyoni et al. [49] explored the adversarial models of traffic analysis attacks on Tor in 2020. The authors examine traffic analysis attacks based on the adopted adversary model and its alignment with Tor’s threat model. However, the paper reviews only ten traffic analysis attacks from the perspective of the adopted adversary model, examining the effectiveness of each within the context of the adopted adversary model and Tor’s threat assumptions. Moreover, this paper lacks an extensive introduction to the WF attack, thus failing to provide a comprehensive landscape of such attacks. Later, Karunanayake et al. [50] surveyed de-anonymization attacks, categorizing them based on the components involved, including entry and exit routers, onion proxy, side channels, and hybrid approaches. However, this survey paper classifies WF attacks under side channels, a categorization that lacks standardization and fails to capture the nature of these attacks accurately. This classification may create ambiguity in distinguishing WF attacks from other de-anonymization techniques. Aminuddin et al. [27] provided a comprehensive survey of WF attacks on the Tor network in 2023. This paper systematically reviews existing WF techniques and categorizes them based on five primary aspects: threat model, victim target, website realm, traffic feature, and traffic classifier. This work also critically examines nine key assumptions that limit the real-world applicability of WF techniques, such as closed-world settings, static content, and passive webpage assumptions. However, the survey focuses almost exclusively on attack methodologies, leaving website fingerprinting defenses outside its scope. Moreover, its discussion of datasets is relatively limited, although dataset assumptions and realism are crucial factors that affect the performance of WF attacks. A deeper integration of dataset considerations with attack and defense methodologies would have provided a more balanced perspective.
@@
Liu et al. [25] presented a survey on WF attacks and defenses, with a primary focus on deep learning (DL) paradigms. Their work reviews the architecture, efficacy, and computational overhead of DL-based models in WF security. However, their study centers predominantly on DL methodologies rather than providing a systematic analysis of attack and defense strategies themselves. Additionally, it lacks deeper insights into the characteristics of WF attacks and defenses, and does not provide an extensive discussion on ongoing challenges, future directions, or evolutionary trends. Chao et al. [28] presented a comprehensive survey of anonymity networks with a particular focus on Tor, I2P, and Freenet. This paper systematically categorizes the vulnerabilities, attack strategies, and defense mechanisms that shape the ongoing battle between anonymity systems and adversaries. However, the survey primarily emphasizes protocol-level vulnerabilities, attacks, and defenses. While the authors briefly discuss some WF attack and defense studies, WF security is not the central focus. Consequently, the paper provides limited insights into WF attacks and defenses, particularly their broader implications for Tor between the client and the Tor entry node. To summarize, there is an urgent need to provide a comprehensive overview of the latest advancements in WF attack techniques or defensive mechanisms for advancing future research.
@@ III-A Tor Architecture
Tor routes traffic through a series of randomly selected intermediary nodes called relays [51, 52, 1]. As shown in Figure 2, when a client initiates a connection, it selects three relays to form a circuit: an entry guard, a middle relay, and an exit relay. The client typically uses the circuit for approximately ten minutes before establishing a new one. The entry guard is the only relay that knows the client’s IP address. The middle relay helps obscure the path between the entry and exit points, and the exit relay is responsible for forwarding the traffic to the final destination. Each relay in the circuit only knows its immediate predecessor and successor, ensuring that no single relay has complete knowledge of both the source and destination, thus preserving user anonymity. When a user (client) initiates a connection through the Tor network using the Tor Browser, the traffic is first encrypted multiple times using a process called onion routing. This layered encryption ensures that no single relay can simultaneously identify both the origin and destination of the communication, reinforcing Tor’s core principle of anonymity.
Tor applies a layered encapsulation and segmentation process to ensure secure and reliable network packet delivery. Specifically, Figure 3 illustrates the data representation across different layers of transport within the Tor network. At the application layer, Tor encapsulates data into fixed-size 512-byte cells, which serve as the fundamental unit of communication. These cells are subsequently encrypted and aggregated into TLS records, with each record typically containing several complete Tor cells. As TLS records traverse the network stack, they are further segmented into multiple TCP packets, constrained by the maximum segment size (MSS) of 1448 bytes for Ethernet. In the Tor network, each circuit can carry multiple streams, with each stream corresponding to a separate TCP connection. To manage flow control, Tor employs SENDME cells: one is sent for every 50 incoming cells per stream and one for every 100 incoming cells per circuit. In this way, Tor can produce traffic patterns that reflect the substantial data volumes generated by modern websites [11, 29].
@@
As shown in TABLE II, Tor traffic protocol (Onion Routing) differs from traditional Hypertext Transfer Protocol Secure (HTTPS) protocols by transmitting data in fixed-size 512-byte cells, which helps obscure traffic patterns and resist analysis. HTTPS is a secure communication protocol that combines HTTP with Transport Layer Security (TLS). It encrypts data between a client (e.g., web browser) and a server, ensuring confidentiality, integrity, and authentication. Unlike HTTPS, QUIC is a protocol that encrypts both application data and transport-layer headers over the User Datagram Protocol (UDP). This protocol modifies the traditional TLS handshake to fit within its streamlined architecture. However, neither HTTPS nor QUIC conceals the client’s IP address, packet size, or routing path, making them less effective against anonymity threats compared to Virtual Private Networks (VPNs) and the Tor network. In contrast, VPN encrypts a user’s internet traffic and routes it through a remote server, masking the user’s IP address and enabling access to geo-restricted content while maintaining relatively fast speeds. However, since the VPN provider can potentially monitor a user’s activity, it is essential to choose a trustworthy service.
@@ III-B Tor Traffic Features
Traffic timing Tor cells are sent and received over time between the client and entry node. Even though the traffic is encrypted, its timing patterns are strongly influenced by a website’s structure, resources, and delivery behavior. Packet inter-arrival time [37]. Random/adaptive dummy packets to vary the timing patterns [53, 54, 55]. Traffic regularization to normalize the timing patterns [56, 57, 58]. Adversarial perturbation to disturb the timing patterns [59, 60, 61].
Time gap [62].
Transmission time/Total trace time [58, 9, 63, 64].
The delay between the first outgoing packet and the first incoming packet [62, 65].
Packet length Different websites produce distinctive sequences of packet lengths because of variations in their webpage payloads. Total number of packets [9, 66]. In real-world scenarios, attackers cannot capture the total traffic length since Tor’s start and end cells are unknown. So, WF defense studies exclude this feature.
Number of outgoing packets [35].
Number of incoming packets [35].
The total length in bytes of the packet [66, 37].
@@
Total packet/cell size [9, 7, 35].
Packet ordering Different websites exhibit distinct request–response patterns, such as the alternation of client requests and server responses, the clustering of consecutive packets in one direction, or the timing of specific sequences. Packet concentration and ordering features list [7, 35]. Add randomness to the traffic stream, such as injecting dummy packets [55, 67, 68] and introducing fake delays [58, 69]. Reordering packets by Traffic morphing [22, 23, 19] and creating a single ”supersequence” of packets [9].
Average of the outgoing packet ordering list [64].
Standard deviation of the outgoing packet ordering list [35].
Time sequence/time-ordered manner [33, 70].
Number of incoming and outgoing cells before and after the candidate cell [37].
Traffic Bursts Traffic bursts are short, high-rate sequences of packets (often dominated by either uploads or downloads) separated by intervals of relatively low activity, arising as a browser fetches different resources such as images, videos, and scripts at various times. The frequency distribution of the packets or cells for a time interval [8, 9, 33, 71]. Obfuscate characteristic packet sequence [16, 11, 72]. Create a consistent flow [58, 57, 73]. Obfuscate traffic flow classification [23, 19, 74, 46].
A sequence of outgoing packets, which is triggered by one incoming packet [11, 36].
Direction The incoming/outgoing (upload/download) traffic flow. Assigning +1 to outgoing packets and -1 to incoming packets [12, 39, 40, 33]. Defense models cannot change the packet direction.
Coarse features In website fingerprinting, coarse features are high-level traffic statistics that reflect distinct patterns in website content and loading behavior, such as bandwidth consumption and time-window-based counting. Total transmission time or total timespan of a trace [75]. Adjusting transmission rates or volumes to flatten bandwidth and hide burstiness. Morphing traffic distributions (e.g., timespan, bandwidth) to mimic other sites.
Total per-direction bandwidth [58].
Traffic ”burstiness” and burst bandwidth [76].
Time-window-based counting [29].

### v2

@@ I Introduction
The Tor network is a decentralized, volunteer-operated system designed to provide website browsing anonymity. Tor applies sophisticated onion routing through a series of relay nodes and encryption mechanisms to anonymize traffic. Its architecture is built to protect users from surveillance, traffic analysis, and censorship by hiding their identity and online activity. Thus, Tor makes adversaries difficult to trace the origin or destination of the data [dingledine2004tor, mani2018understanding, dantas2020detection, panchenko2011website].
However, Tor’s network traffic can still be susceptible to traffic analysis techniques, including website fingerprinting (WF) attacks [cherubin2022online, juarez2014critical, cai2012touching]. By exploiting patterns in encrypted traffic, these attacks aim to expose the websites a user is visiting. Distinctive patterns in packet size [herrmann2009website, wang2014effective, abe2016fingerprinting], timing [panchenko2016website, rahman2019tik, yin2024traces], and flow direction [rahman2019tik, yan2018feature] can still leak identifying information about the websites being accessed. In response, a variety of defense mechanisms [cai2014systematic, juarez2016toward, nithyanand2014glove] have been developed to obscure or randomize traffic patterns, thereby reducing the risk of identification. These strategies include traffic obfuscation [liu2023smart, shen2024real], adversarial perturbation [hou2021universal, sun2022practical], and website morphing [chan2018website, al2019bimorphing, ling2024wfguard]. While effective in mitigating fingerprinting attempts, such defenses must carefully balance privacy protection with performance to ensure they do not degrade the user experience.
@@ I-A Motivation
Despite the growing significance of WF attack and defense approaches, a systematic and comprehensive survey of the field remains lacking in academia. Existing surveys primarily focus on WF attacks and defenses using machine learning (ML) models [liu2023survey, shen2022machine, aminuddin2023rise]. However, these works provide limited insights into the design of models in critical aspects such as concept drift, single-tab versus multi-tab browsing scenarios, and dynamic adversarial capabilities. Moreover, while previous surveys [aminuddin2023rise, chao2024systematic] introduce several defense mechanisms, their taxonomies are neither systematic nor sufficiently granular. For instance, Chao et al. [chao2024systematic] categorize WF attacks and defenses within anonymity networks based on different protocol layers of the onion route. However, a gap remains in understanding the comparative effectiveness of various techniques under diverse threat models and defense scenarios. Existing work fails to adequately categorize defenses according to design principles (e.g., obfuscation, padding, traffic morphing) or threat model alignment, limiting their utility for comparative evaluation. Furthermore, datasets in shaping WF research are not well justified. These surveys fail to categorize defenses based on factors such as design principle (e.g., obfuscation, padding, traffic morphing) or threat model alignment adequately. For example, the existing survey [aminuddin2023rise] discusses dataset features and assumptions, but does not examine the availability, diversity, and limitations of existing datasets in depth. Therefore, a systematic review of WF attacks and defenses in Tor is imperative. Unlike existing surveys, our work provides an in-depth review of available datasets and examines their features to guide dataset selection for future research. Building on this foundation, we present a comprehensive review of WF attack approaches and categorize defense strategies according to their underlying mechanisms, highlighting their respective strengths and weaknesses against state-of-the-art (SOTA) attack models.
@@ I-B1 Datasets
Labeled datasets mapping traffic traces to websites are required in the SOTA WF attacks [deng2025countmamba, mathews2024laserbeak, gao2024multi, deng2024robust, zhao2024towards], which usually employ supervised learning techniques. The effectiveness of existing ML models largely depends on the quality, diversity, and scale of the datasets used. Therefore, the development of realistic, diverse, and up-to-date datasets is essential for advancing research in this field and ensuring the reliability of experimental results [zhu2024toscorr].
There are two types of datasets: closed-world and open-world. A closed-world dataset assumes that the user visits only a predefined set of websites (e.g., the top 100 sites from Alexa), and both the attacker and defender know this set in advance [wang2013improved]. The closed-world dataset is better suited for studying the upper bound of attack performance in a controlled setting, as it facilitates comparisons between different WF models under ideal conditions. However, closed-world evaluation overlooks the fact that users can visit websites outside the training set of an ML model. In contrast, an open-world dataset contains a smaller set of monitored websites (the attacker is interested in) and a much larger set of unmonitored websites representing background traffic. In this setting, the attacker must distinguish between monitored and unmonitored sites. Open-world datasets more accurately reflect real-world usage, allowing researchers to evaluate the effectiveness of attacks and defenses under more realistic assumptions [wang2014effective, hayes2016k].
@@ I-B2 WF attacks
The WF attacks section introduces the existing website fingerprinting attacks by providing a comprehensive overview of the various techniques that have been developed over time. This includes a detailed discussion of the underlying models used, ranging from traditional ML algorithms such as k-Nearest Neighbors (k-NN) [wang2014effective, zhao2024towards, wang2013improved] and Support Vector Machines (SVM) [wang2013improved, wang2016realistically, panchenko2011website] to more advanced DL approaches such as Convolutional Neural Networks (CNN) [hong2024website, shi2025multiscale, zou2024relation, yin2024traces, chawla2024espresso, xu2025apwf] and Transformer-based architectures [jin2023transformer, mathews2024laserbeak, ding2024multi]. The section also covers the process of feature selection, highlighting which traffic characteristics, such as packet sizes, directions, timing, and burst patterns, are most informative for classification. Additionally, the performance of different attacks is compared using standard evaluation metrics, such as accuracy, precision, and recall, across both closed-world and open-world scenarios. This comparison can help illustrate the evolution of WF attacks and provide insight into which methods are most effective under various conditions.
@@ I-B3 WF defenses
The WF defenses section introduces the existing techniques designed to counter website fingerprinting attacks, with a particular focus on a range of key defense mechanisms. Among these, adaptive padding is a dynamic approach that injects dummy traffic based on live traffic patterns to conceal burst and packet size features, striking a balance between privacy and bandwidth overhead. Traffic morphing attempts to transform the traffic of one website to resemble that of another, thereby confusing fingerprinting classifiers by altering key statistical features [wang2014effective, wang2017walkie, al2019bimorphing, ling2024wfguard]. Adversarial perturbation techniques borrow principles from adversarial ML to proactively disrupt attack models [gong2022surakav, qiao2024trace, jiang2024rudolf], where minor, intentional modifications to network traffic are used to mislead ML models. Each method is discussed in terms of its effectiveness, overhead, implementation complexity, and suitability for real-world Tor systems. This in-depth review will help readers understand the current state of WF defenses and the challenges that remain in developing robust, low-cost, and scalable solutions.
@@ II Related Works
Basyoni et al. [basyoni2020traffic] 2020 Traffic Analysis Attacks on Tor.
Karunanayake et al. [karunanayake2021anonymisation] 2021 De-Anonymisation Attacks on Tor.
Aminuddin et al. [aminuddin2023rise] 2023 Analysis on Tor-based WF techniques and assumptions.
Liu et al. [liu2023survey] 2023 Deep learning on Tor-based WF attacks and defenses.
Chao et al. [chao2024systematic] 2024
@@
Basyoni et al. [basyoni2020traffic] explored the adversarial models of traffic analysis attacks on Tor in 2020. The authors examine traffic analysis attacks based on the adopted adversary model and its alignment with Tor’s threat model. However, the paper reviews only ten traffic analysis attacks from the perspective of the adopted adversary model, examining the effectiveness of each within the context of the adopted adversary model and Tor’s threat assumptions. Moreover, this paper lacks an extensive introduction to the WF attack, thus failing to provide a comprehensive landscape of such attacks. Later, Karunanayake et al. [karunanayake2021anonymisation] surveyed de-anonymization attacks, categorizing them based on the components involved, including entry and exit routers, onion proxy, side channels, and hybrid approaches. However, this survey paper classifies WF attacks under side channels, a categorization that lacks standardization and fails to capture the nature of these attacks accurately. This classification may create ambiguity in distinguishing WF attacks from other de-anonymization techniques. Aminuddin et al. [aminuddin2023rise] provided a comprehensive survey of WF attacks on the Tor network in 2023. This paper systematically reviews existing WF techniques and categorizes them based on five primary aspects: threat model, victim target, website realm, traffic feature, and traffic classifier. This work also critically examines nine key assumptions that limit the real-world applicability of WF techniques, such as closed-world settings, static content, and passive webpage assumptions. However, the survey focuses almost exclusively on attack methodologies, leaving website fingerprinting defenses outside its scope. Moreover, its discussion of datasets is relatively limited, although dataset assumptions and realism are crucial factors that affect the performance of WF attacks. A deeper integration of dataset considerations with attack and defense methodologies would have provided a more balanced perspective.
@@
Liu et al. [liu2023survey] presented a survey on WF attacks and defenses, with a primary focus on deep learning (DL) paradigms. Their work reviews the architecture, efficacy, and computational overhead of DL-based models in WF security. However, their study centers predominantly on DL methodologies rather than providing a systematic analysis of attack and defense strategies themselves. Additionally, it lacks deeper insights into the characteristics of WF attacks and defenses, and does not provide an extensive discussion on ongoing challenges, future directions, or evolutionary trends. Chao et al. [chao2024systematic] presented a comprehensive survey of anonymity networks with a particular focus on Tor, I2P, and Freenet. This paper systematically categorizes the vulnerabilities, attack strategies, and defense mechanisms that shape the ongoing battle between anonymity systems and adversaries. However, the survey primarily emphasizes protocol-level vulnerabilities, attacks, and defenses. While the authors briefly discuss some WF attack and defense studies, WF security is not the central focus. Consequently, the paper provides limited insights into WF attacks and defenses, particularly their broader implications for Tor between the client and the Tor entry node. To summarize, there is an urgent need to provide a comprehensive overview of the latest advancements in WF attack techniques or defensive mechanisms for advancing future research.
@@ III-A Tor Architecture
Tor routes traffic through a series of randomly selected intermediary nodes called relays [mccoy2008shining, snader2008tune, dingledine2004tor]. As shown in Figure 2, when a client initiates a connection, it selects three relays to form a circuit: an entry guard, a middle relay, and an exit relay. The client typically uses the circuit for approximately ten minutes before establishing a new one. The entry guard is the only relay that knows the client’s IP address. The middle relay helps obscure the path between the entry and exit points, and the exit relay is responsible for forwarding the traffic to the final destination. Each relay in the circuit only knows its immediate predecessor and successor, ensuring that no single relay has complete knowledge of both the source and destination, thus preserving user anonymity. When a user (client) initiates a connection through the Tor network using the Tor Browser, the traffic is first encrypted multiple times using a process called onion routing. This layered encryption ensures that no single relay can simultaneously identify both the origin and destination of the communication, reinforcing Tor’s core principle of anonymity.
Tor applies a layered encapsulation and segmentation process to ensure secure and reliable network packet delivery. Specifically, Figure 3 illustrates the data representation across different layers of transport within the Tor network. At the application layer, Tor encapsulates data into fixed-size 512- or 514-byte cells, which serve as the fundamental unit of communication. These cells are subsequently encrypted and aggregated into TLS records, with each record typically containing several complete Tor cells. As TLS records traverse the network stack, they are further segmented into multiple TCP packets, constrained by the maximum segment size (MSS) of 1448 bytes for Ethernet. In the Tor network, each circuit can carry multiple streams, with each stream corresponding to a separate TCP connection. To manage flow control, Tor employs SENDME cells: one is sent for every 50 incoming cells per stream and one for every 100 incoming cells per circuit. In this way, Tor can produce traffic patterns that reflect the substantial data volumes generated by modern websites [panchenko2016website, deng2025countmamba].
@@
As shown in TABLE II, Tor traffic protocol (Onion Routing) differs from traditional Hypertext Transfer Protocol Secure (HTTPS) protocols by transmitting data in fixed-size 512- or 514-byte cells, which helps obscure traffic patterns and resist analysis. HTTPS is a secure communication protocol that combines HTTP with Transport Layer Security (TLS). It encrypts data between a client (e.g., web browser) and a server, ensuring confidentiality, integrity, and authentication. Unlike HTTPS, QUIC is a protocol that encrypts both application data and transport-layer headers over the User Datagram Protocol (UDP). This protocol modifies the traditional TLS handshake to fit within its streamlined architecture. However, neither HTTPS nor QUIC conceals the client’s IP address, packet size, or routing path, making them less effective against anonymity threats compared to Virtual Private Networks (VPNs) and the Tor network. In contrast, VPN encrypts a user’s internet traffic and routes it through a remote server, masking the user’s IP address and enabling access to geo-restricted content while maintaining relatively fast speeds. However, since the VPN provider can potentially monitor a user’s activity, it is essential to choose a trustworthy service.
@@ III-B Tor Traffic Features
Traffic timing Tor cells are sent and received over time between the client and entry node. Even though the traffic is encrypted, its timing patterns are strongly influenced by a website’s structure, resources, and delivery behavior. Packet inter-arrival time [wang2016realistically]. Random/adaptive dummy packets to vary the timing patterns [shmatikov2006timing, cui2018realistic, pulls2020towards]. Traffic regularization to normalize the timing patterns [lu2018dynaflow, gong2020zero, dyer2012peek]. Adversarial perturbation to disturb the timing patterns [hou2020wf, rahman2020mockingbird, nasr2021defeating].
Time gap [xu2018multi].
Transmission time/Total trace time [dyer2012peek, wang2014effective, kwon2015circuit, attarian2019adawfpa].
The delay between the first outgoing packet and the first incoming packet [xu2018multi, yin2021automated].
Packet length Different websites produce distinctive sequences of packet lengths because of variations in their webpage payloads. Total number of packets [wang2014effective, liberatore2006inferring]. In real-world scenarios, attackers cannot capture the total traffic length since Tor’s start and end cells are unknown. So, WF defense studies exclude this feature.
Number of outgoing packets [wang2013improved].
Number of incoming packets [wang2013improved].
The total length in bytes of the packet [liberatore2006inferring, wang2016realistically].
@@
Total packet/cell size [wang2014effective, cai2012touching, wang2013improved].
Packet ordering Different websites exhibit distinct request–response patterns, such as the alternation of client requests and server responses, the clustering of consecutive packets in one direction, or the timing of specific sequences. Packet concentration and ordering features list [cai2012touching, wang2013improved]. Add randomness to the traffic stream, such as injecting dummy packets [pulls2020towards, abusnaina2020dfd, kadianakis2021tor] and introducing fake delays [dyer2012peek, cai2014cs]. Reordering packets by Traffic morphing [chan2018website, al2019bimorphing, shen2024real] and creating a single ”supersequence” of packets [wang2014effective].
Average of the outgoing packet ordering list [attarian2019adawfpa].
Standard deviation of the outgoing packet ordering list [wang2013improved].
Time sequence/time-ordered manner [zhao2024towards, zhou2023wf].
Number of incoming and outgoing cells before and after the candidate cell [wang2016realistically].
Traffic Bursts Traffic bursts are short, high-rate sequences of packets (often dominated by either uploads or downloads) separated by intervals of relatively low activity, arising as a browser fetches different resources such as images, videos, and scripts at various times. The frequency distribution of the packets or cells for a time interval [herrmann2009website, wang2014effective, zhao2024towards, bahramali2023realistic]. Obfuscate characteristic packet sequence [juarez2016toward, panchenko2016website, lu2023lightweight]. Create a consistent flow [dyer2012peek, gong2020zero, holland2020regulator]. Obfuscate traffic flow classification [al2019bimorphing, shen2024real, li2022minipatch, gong2022surakav].
A sequence of outgoing packets, which is triggered by one incoming packet [panchenko2016website, hayes2016k].
Direction The incoming/outgoing (upload/download) traffic flow. Assigning +1 to outgoing packets and -1 to incoming packets [rahman2019tik, shi2025multiscale, zou2024relation, zhao2024towards]. Defense models cannot change the packet direction.
Coarse features In website fingerprinting, coarse features are high-level traffic statistics that reflect distinct patterns in website content and loading behavior, such as bandwidth consumption and time-window-based counting. Total transmission time or total timespan of a trace [tan2024inter]. Adjusting transmission rates or volumes to flatten bandwidth and hide burstiness. Morphing traffic distributions (e.g., timespan, bandwidth) to mimic other sites.
Total per-direction bandwidth [dyer2012peek].
Traffic ”burstiness” and burst bandwidth [zhu2025wf].
Time-window-based counting [deng2025countmamba].